OpenStudy (anonymous):

Simplify 1.25 into a percentage.

Directrix (directrix):

% is hundredths. 1 = 100/100 = 100% ,25 = 25/100 = 25 % 1.25 = 125%

Directrix (directrix):

If you want to review the concept, go here: http://www.coolmath.com/prealgebra/03-percents/04-percents-convert-decimal-to-percent-01